(1) The following thermochemical equation is for the reaction of ethylene(g) with water(g) to form ethanol(g) (CH3CH2OH).

C2H4(g) + H2O(g) CH3CH2OH(g) H = -45.6 kJ
How many grams of C2H4(g) would have to react to produce 10.8 kJ of energy?

(2) The following thermochemical equation is for the reaction of hydrogen peroxide(l) to form water(l) and oxygen(g).

2H2O2(l)2H2O(l) + O2(g) H = -196 kJ
How many grams of H2O2(l) would have to react to produce 28.5 kJ of energy?

(3) The following thermochemical equation is for the reaction of carbon dioxide(g) with hydrogen(g) to form carbon monoxide(g) and water(g).

CO2(g) + H2(g) CO(g) + H2O(g) H = 41.2 kJ
When 14.5 grams of carbon dioxide(g) react with excess hydrogen(g),____ kJ of energy are____ .
1
Expert's answer
2020-04-27T01:45:14-0400

1."45.6 kJ" of energy is produced by 1 mole of C2H4(g).

Thus, "10.8 kJ" of energy is produced by "10.8\/45.6=0.237" mol of C2H4(g)"=0.237*28=6.632 gms."


2."196 kJ" of energy is produced by 2 moles of H2O2(l).

Thus, "28.5 kJ" of energy is produced by "2*28.5\/196=0.291" mol of H2O2(l)"=0.291*34=9.888 gms."


3."41.2 kJ" of energy is produced by 1 mole of CO2(g) or 44gm of CO2.

Thus, "44gm" of CO2 produces "41.2 kJ" of energy.

Thus, "14.5gm" of CO2 produces "14.5*41.2\/44 =13.577 kJ" of energy.